OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8. Show all work.

OpenStudy (anonymous):

y=k/3x 4=k/3*6 k=4*3*6=72 Then 8=72/3x 8=24/x x=24/8=3